OpenStudy (anonymous):

My teacher admitted that my book didn't explain how to do this well enough..I need help..Any ideas? Award At End. Given r(t), find the following: r(t) = < t^2, sin(t) - tcos(t), cos(t) + tsin(t) > Write a(t) in terms of its tangential and normal components.

OpenStudy (dan815):

x=t^2, y=sint-tcost, z=cost + tsint

OpenStudy (dan815):

a(t)= k N(t) + b (T(t)) lets assume a(T) can be written as a linear combination of these 2 vectors

OpenStudy (dan815):

lets find the tangetial component that is not too hard

OpenStudy (dan815):

wait!

OpenStudy (dan815):

i gotta refresh its been a really long time ive derived these

OpenStudy (dan815):

do you know how to find the unit tangent vetors first

OpenStudy (anonymous):

Isn't it v(t) T(t) = ||v(t)||?

OpenStudy (anonymous):

Wow I can not work the equation opetator. @dan815 xD

OpenStudy (anonymous):

Basically T(t) = v(t)/||v(t)||

OpenStudy (anonymous):

...help...